A 25-year, 8% semiannual coupon bond with a par value of $1,000 may be called in 4 years at a call price of $1,100. The bond sells for $950. (Assume that the bond has just been issued.) a. What is the bonds yield to maturity? Round your answer to two decimal places b. What is the bonds current yield? Round your answer to two decimal places. c. What is the bonds capital galn or loss yleld? Loss should be Indicated with minus sign. Round your answer to two declmal places. d. What is the bonds yleld to call? Round your answer to two decimal places
Answer #1

a. Yield to maturity = (I+((F-P)/n))/((F+P)/2), where:

  • I = periodic coupon payment
  • F=Face value
  • P=current price
  • n=number of periods over which bond will be repaid

YTM =((0.08/2*1000)+((1000-950)/(25*2)))/((1000+950)/2) = 4.24%, this is a semi-annual yield.

Annual YTM = 4.24*2 = 8.48%

b. Current yield = yearly coupon payment/Current market price

=(0.08*1000)/950 = 8.42%

c. Capital gain yield on bond = YTM-Current yield = 8.48-8.42 = 0.06%

d. Yield to Call = (I+((F-C)/t))/((F+C)/2),where:

  • I = periodic coupon payment
  • F= Face value
  • C=call price
  • t=periods till call date

YTC=((0.08/2*1000)+((1100-950)/(4*2)))/((1100+950)/2) = 5.76%, this is semi-annual yield

Annualized YTM = 5.76%*2 = 11.52% (approx)